Under Lockdown restrictions after EFB confirmed two weeks ago.   Weather  20 degrees, cloudy.
Hive#1  Inspected by David Parkinson.   Brood Chamber: 14 x 12.
QNS,   Lots of Qc’s after killing Q last week, all taken down bar two best ones on frame 7.              6 f brood,   3 f stores,  room ok,  1 super most empty.  Temper 8.
Hive#2  Inspected by Liz.    Brood Chamber: 14 x 12.
QS,  8 play cups ,not charged,  8 f brood, stores and room OK. 1 super filling. Temper 7,   Not yet pulling shallow varroa trap.  3 stings, not happy bees.
Hive#4  Inspected by Gaye.    Brood chamber:  National Deep.
QNS,   1 torn down Qc,  8 f brood BAS,  stores and room ok,  3 supers, 2 full.  Temper 8.   Swapped  some of the full super frames for empty with Hive#1.
Hive#5  Inspected by David,   Brood Chamber:  Deep and  a Half.
QNS,  6 Qc’s,   4 f sealed brood, no eggs, some larva, stores and room ok.  2 supers. Removed all but 2 Qc’s as poss queenless.  1 frame with 2 Qc’s given to Hive#8.   Very aggressive bees, Temper 4.  Lots of stings.( poss due to being Queenless)
Hive#6   Inspected by Richard,  Brood Chamber: Deep and a half.
QS,  No Qc’s, 8d f brood,  6s f brood all stages, stores and room ok,  1 super nearly full. Temper 9
Hive#7   Left alone as poss Virgin queen,   ( Qc’s left 22/5)  check next week.
Hive#7A  Not inspected.
Hive#8  Inspected by Richard.   Brood Chamber: Deep.
No Queen,  No Qc’s.  Little brood mainly Drone left. Lots of stores and room.  Added frame of brood with 2 Qc’s, 1 open 1 sealed, from Hive#5 ( no bees).   Bees very agitated. Temper 6.
Hive#10  Inspected by Liz,    Top Bar Hive.
QNS,  no Qc’s,  3 f BAS.   2 f stores, room ok.   Temper 10,   Wax moth still evident and ants a problem.
Hive#15F   Inspected by Gaye,  Nuc in National Deep.
QNS, no Qc’s,  6 f Brood . BAS.  1.5 f stores, room ok, No super.   Temper 10.
Hive#15H  Inspected by Liz.    Nuc in Deep.
QNS, no Qc’s,   5 f BAS and 1 day old eggs,  stores and room ok,  1 super.   Moved bees 3 ft towards  no 3 position,  before inspection, so bees very agitated, better to move after inspection.
Nuc’s   16 A and 16 B   (both made up from colony 4) Not gone through but bees flying ok.